MARC:

If (x-a) is a factor of the expression ax^3-3x^2-5ax-9,find the possible values of a.Factorise the expression for each of the values of a.

MARC:

@Zepdrix

Zepdrix:

If (x-a) is a factor of the polynomial, then x=a is a root of the polynomial. Plugging in a should give us zero, \(\large\rm a\cdot a^3-3a^2-5a(a)-9=0\)

MARC:

I tried using the method u taught me before. ^

Zepdrix:

Does this end up being some weird thing? Hmm let's see..

Zepdrix:

\(\large\rm a^4-3a^2-5a^2-9=0\) Combining like-terms, \(\large\rm a^4-8a^2-9=0\)

Zepdrix:

Hmm this one works out pretty nicely I think. -9 = -9*1 and -8 = -9+1

Zepdrix:

So ya, we'd have to apply that little trick from last time, \(\large\rm u=a^2\)

MARC:

yep,I got \((a^2-9)(a^2+1)\)

Zepdrix:

\(\large\rm a^4-8a^2-9=0\) Ok so we've done this so far, \(\large\rm (a^2-9)(a^2+8)=0\) don't forget about conjugates! :) \(\large\rm (a-3)(a+3)(a^2+8)=0\)

Zepdrix:

So that takes care of the second part, factorize completely. To find values for a, we can apply our Zero-Factor Property, setting each individual factor equal to zero, \(\large\rm (a-3)=0\) \(\large\rm (a+3)=0\) \(\large\rm (a^2+8)=0\)

Zepdrix:

And solve for a in each case.

MARC:

I thought it should be \((a^2-9)(a^2+\color{red}{1})\)

Zepdrix:

Ahhh sorry >.<

Zepdrix:

Did you understand the conjugates part though? How we got to (a-3)(a+3) ?

MARC:

It's okay ...yes,I understand. I managed to factorise the expression \((a-3)(a+3)(a+1)(a-1)\)

Zepdrix:

No no, sum of squares does not factor like that. \(\large\rm a^2+1\ne(a+1)(a-1)\)

MARC:

is it \(\(a+1)^2\) ?

Zepdrix:

\(\large\rm a^2+1=a^2+1\) Can't do anything nice with it. if you apply your zero-factor property, then you might see why. \(\large\rm a^2+1=0\) \(\large\rm a^2=-1\) \(\large\rm a=\pm\sqrt{-1}\) This gives us no real solutions right?

MARC:

oh yesss... ^

Zepdrix:

So then this is our full factorization, \(\large\rm (a-3)(a+3)(a^2+1)\)

MARC:

my answer in the book is a little bit different. xD answer: a=3,\(3(x-3)(x+1)^2\),a=-3,\(-3(x+3)(x-1)^2\)

Zepdrix:

whu 0_o

Zepdrix:

Ohhh I see. My bad. I completely forgot about the expression involving x.

Zepdrix:

\(\large\rm (a-3)(a+3)(a^2+1)\) From that factorization, we conclude that the values of a are 3 and -3. So now we need to plug these values in for our a's, and factorize that expression, \(\large\rm \color{indianred}{a}x^3-3x^2-5\color{indianred}{a}x-9\) So for a=3, \(\large\rm \color{indianred}{3}x^3-3x^2-15x-9\)

MARC:

okay

Zepdrix:

Hmm we have an ugly cubic so we can umm...

Zepdrix:

Oh, well we already know that (x-a) is a factor of this polynomial. So in this case, (x-3) is a factor. So we can do long division or synthetic.

Zepdrix:

This problem is pretty brutal, really long lol

MARC:

Yep xD

MARC:

Can I use calculator for now? xD

Zepdrix:

ya

MARC:

Okay,it will be easier. ^

MARC:

I got (x-3)(x+1)

MARC:

Hmm,something is wrong here. My calculator is showing two solution

Zepdrix:

We can start by factoring a 3 out of everything, \(\large\rm 3(x^3-x^2-5x-3\) And then pulling the (x-3) factor out gives us, \(\large\rm 3(x-3)(x^2+2x+1)\) Which factors into, \(\large\rm 3(x-3)(x+1)^2\)

MARC:

Then,we do the samething for a=-3

Zepdrix:

Ya i guess so XD

MARC:

okay,I think I know how to do. Thank you! @Zepdrix

Zepdrix:

cool :)